# Turb-L1: Achieving Long-term Turbulence Tracing By Tackling Spectral Bias

### v2

@@ 2.4 Hierarchical Dynamics Synthesis for Full-Spectrum Fidelity
The Hierarchical Dynamics Synthesis (HDS) mechanism, $\mathcal{HDS}(\cdot;\theta_{\mathcal{HDS}})$, is pivotal to Turb-L1’s strategy for mitigating spectral bias. It operates on the latent representation $\mathbf{z}_{\text{enc}}\in\mathcal{M}_{\omega}$ from the multi-grid encoder $\mathcal{E}$, precisely evolving turbulence dynamics within this manifold. The HDS design leverages insights into the distinct frequency responses of neural operators: Convolutional Neural Networks (CNNs) excel at capturing high-frequency local details, whereas Transformer-like architectures are adept at integrating low-frequency global information park2022vision ; schiaffini2024neural . Turb-L1 architecturally synergizes these complementary properties for precise, high-fidelity, full-spectrum turbulence evolution.
@@
For synthesizing high-frequency dynamic primitives $\mathbf{h}_{\omega_{H}}$, $\mathcal{T}_{\omega_{H}}$ employs operators highly sensitive to local variations and high-frequency content. Convolutional operations, particularly with small receptive field kernels $\kappa_{j}^{(H)}$, inherently function as high-pass filters, adept at extracting high-frequency signals such as edges and textures. Consequently, $\mathcal{T}_{\omega_{H}}$ is primarily instantiated via such convolutional architectures:
[…]
where $\sigma$ is an activation and $\mathbf{b}_{j}^{(H)}$ a bias. This design is aimed at enhancing gradient-like information and fine vortex features within $\mathbf{z}_{\text{enc}}$.
Conversely, synthesizing low-frequency dynamic primitives $\mathbf{h}_{\omega_{L}}$ is achieved by $\mathcal{T}_{\omega_{L}}$, which targets large-scale structures and global flow trends. The self-attention mechanism (MSA) in Transformers, by performing a weighted average across all input elements, inherently acts as a low-pass filter, effectively integrating global context. $\mathcal{T}_{\omega_{L}}$ is thus instantiable as a Global Adaptive Weighted Aggregation (GAWA) mechanism, rooted in MSA:
$[\mathcal{T}_{\omega_{L}}(\mathbf{z}_{\text{enc}})]_{i}=\sum_{k=1}^{N_{h}}\mathbf{W}_{k}^{(O)}\left(\sum_{j=1}^{N_{T}}\alpha_{ij}^{(k)}(\mathbf{W}_{k}^{(V)}\mathbf{z}_{j})\right),$ (8)
where $\alpha_{ij}^{(k)}=\text{softmax}_{j}\left(\frac{(\mathbf{W}_{k}^{(Q)}\mathbf{z}_{i})^{T}(\mathbf{W}_{k}^{(K)}\mathbf{z}_{j})}{\sqrt{d_{k}}}\right)$, $N_{T}$ is the token count, and $\mathbf{W}_{k}^{(\cdot)}$ are learnable head projection matrices.
The concurrently generated dynamic primitives $\{\mathbf{h}_{\omega_{H}},\mathbf{h}_{\omega_{M}},\mathbf{h}_{\omega_{L}}\}$, conveying distinct frequency channel information, are subsequently fused by a Multi-scale Dynamic Aggregation (MDA) module, $\mathcal{A}_{\text{MDA}}(\cdot,\cdot,\cdot;\theta_{\mathcal{HDS}}^{(\text{agg})})$. Realized as a sequence of convolutional blocks, $\mathcal{A}_{\text{MDA}}$ learns to effectively combine these multi-frequency features, capturing their non-linear dependencies and facilitating cross-frequency interactions. This process yields a unified latent representation $\mathbf{z}_{\text{HDS\_out}}$ encapsulating the full-spectrum evolved dynamics:
@@
As depicted in Figure 2, the HDS output $\mathbf{z}_{\text{HDS\_out}}$ is subsequently channeled to the multi-grid decoder $\mathcal{D}$. The core HDS strategy, involving explicit extraction of frequency-specific dynamics, their dedicated processing, and intelligent aggregation via the convolutional MDA in latent space, forces the model to learn and preserve critical high-frequency details. By accurately capturing global flow while maintaining these fine structures, this design fundamentally mitigates spectral bias, enabling Turb-L1 to trace complex, multi-scale turbulence evolution with high fidelity.
@@ 2.5 Mapping to Physical Space
The Multi-Grid Decoder $\mathcal{D}$ receives the latent space features $\mathbf{z}_{\text{HDS\_out}}$, evolved by the HDS module. Through a sequence of sophisticated hierarchical operations, it aims to reconstruct the turbulent vorticity field $\hat{\omega}(\mathbf{x},t_{k+1})$ in the physical space with maximal fidelity. This process initiates from the coarsest latent representation (corresponding to $\mathbf{z}_{\text{HDS\_out}}$). It iteratively applies a Prolongation Operator $\mathcal{P}_{l}$ (upsampling features from level $l+1$ to $l$) followed by a learnable Feature Refinement Module $\psi_{l}(\cdot;\theta_{\mathcal{D}}^{(l)})$ (refining features at level $l$). This nested sequence of composite operations, where $l$ ranges from $L-1$ down to 00, progressively restores the latent representation. Finally, an output transformation layer $\mathcal{T}_{\text{out}}(\cdot;\theta_{\mathcal{D}}^{(\text{out})})$ maps this fully refined fine-level representation to the final vorticity field prediction. The synergy between the decoder and the HDS mechanism ensures the effective reproduction of full-spectrum dynamical information, particularly high-frequency fine structures, thereby achieving high-fidelity and physically consistent tracking of turbulence evolution. Its overall mathematical formulation is:
$\hat{\omega}(\mathbf{x},t_{k+1})=\mathcal{T}_{\text{out}}\left(\psi_{0}\left(\mathcal{P}_{0}\left(\dots\psi_{L-1}\left(\mathcal{P}_{L-1}\left(\mathbf{z}_{\text{HDS\_out}}\right);\theta_{\mathcal{D}}^{(L-1)}\right)\dots\right);\theta_{\mathcal{D}}^{(0)}\right);\theta_{\mathcal{D}}^{(\text{out})}\right)$ (10)
where $\mathbf{z}_{\text{HDS\_out}}$ is the output from the HDS module, $\mathcal{P}_{l}$ denotes the Prolongation operator at stage $l$, $\psi_{l}(\cdot;\theta_{\mathcal{D}}^{(l)})$ is the Feature Refinement Module with parameters $\theta_{\mathcal{D}}^{(l)}$ for stage $l$ (operating on the output of $\mathcal{P}_{l}$), and $\mathcal{T}_{\text{out}}(\cdot;\theta_{\mathcal{D}}^{(\text{out})})$ is the final output transformation layer with parameters $\theta_{\mathcal{D}}^{(\text{out})}$. The ‘…‘ indicates the continuation of the nested application of $(\psi\circ\mathcal{P})$ blocks for intermediate stages.
We theoretically analyze the link between a model’s spectral accuracy, particularly for high-frequency components, and its long-term autoregressive stability. Consider the true state evolution $\bm{\omega}(t_{k+1})=S(\bm{\omega}(t_{k}))$ and its model approximation $\hat{\bm{\omega}}(t_{k+1})=\mathcal{F}(\hat{\bm{\omega}}(t_{k}))$. The prediction error $\bm{\epsilon}(t_{k})=\hat{\bm{\omega}}(t_{k})-\bm{\omega}^{*}(t_{k})$ evolves approximately as:
@@
Discussion of association with Turb-L1. Theorem A.1 implies that a smaller $\delta_{\text{HF}}$ (i.e., higher model accuracy for high-frequency components) is critical for mitigating error growth in systems sensitive to high-frequency perturbations ($C_{A}\gamma_{\text{high}}>1$). The Turb-L1 architecture, with its Hierarchical Dynamics Synthesis (HDS) mechanism and multi-grid components, is designed to minimize $\delta_{\text{HF}}$ by explicitly addressing Spectral Bias and ensuring high-fidelity information processing across scales. This architectural focus on accurately capturing high-frequency dynamics leads to a significant reduction in the model-induced error term in Eq. (13), thereby enhancing long-term stability in systems like turbulence. This provides a theoretical basis for Turb-L1’s observed performance in maintaining prediction fidelity over extended rollouts. Our experiments also corroborate the above viewpoint, and additional evidence is provided in the Appendix.
@@ 3 Numerical Experiments
Table 1: Performance comparison on isotropic turbulence datasets measured by $L^{2}$ error at different prediction steps. Lower $L^{2}$ error values indicate better performance.
@@
[…]
@@
[…]
@@
[…]
@@ 3.1 Experiment Setup
Datasets  We test Turb-L1 on two classic datasets of two-dimensional isotropic turbulence, a canonical problem in fluid dynamics. These datasets are designed to evaluate the model’s capability to capture complex chaotic dynamics under different physical regimes. ✠ Dataset I: Forced Isotropic Turbulence Benchmark. This scenario follows the Navier-Stokes equation benchmark proposed in Li et al. li2021fourier . The kinematic viscosity is set to $\nu=10^{-5}$. The initial vorticity field, $\omega_{0}$, is sampled from a zero-mean Gaussian random field with a covariance operator $(-\Delta+\tau^{2}I)^{-\alpha/2}$, where $\Delta$ is the Laplacian operator, $I$ is the identity operator, and $\tau$ and $\alpha$ are parameters controlling the spectral shape. The energy density $E(k)$ exhibits a decay law of $E(k)\sim(k^{2}+\tau^{2})^{-\alpha}$. The flow is driven by a fixed, low-wavenumber external force, and no additional drag effects are considered. ✠ Dataset II: Decaying Isotropic Turbulence (McWilliams, 1984). mcwilliams1984emergence This scenario simulates the phenomenon of decaying turbulence. The power spectrum of the initial streamfunction, characterized by $|\hat{\psi}(\mathbf{k})|^{2}\sim k^{-1}(\tau_{0}^{2}+(k/k_{0})^{4})^{-1}$, where $k_{0}$ is the characteristic wavenumber and $\tau_{0}$ is a spectral shape parameter, defines the initial flow field. The initial conditions facilitate a slow energy decay, and the enstrophy density evolves to exhibit features akin to a Kolmogorov energy cascade. The turbulence is unforced and evolves naturally. For more details, see the Appendix.
[…]
@@
Experiment settings  We train all models in this paper on a server equipped with eight NVIDIA A100 GPUs, each with 40GB of memory. We use the DistributedDataParallel mode from the PyTorch paszke2019pytorch framework for distributed training to accelerate convergence. Model inference is performed on a single NVIDIA A100 GPU. The software environment for our experiments is based on Python 3.8. We use PyTorch version 1.8.1 (CUDA 11.1), TorchVision version 0.9.1. During the training process, we consistently use the following hyperparameter configuration: the batch size is set to 20, the total number of epochs is 500, and the initial learning rate is 0.001. To ensure the reproducibility of our experiments, we fix all random seeds to 42. For more details, see the Appendix.

### v3

@@ 3.4 Hierarchical Dynamics Synthesis for Full-Spectrum Fidelity
The Hierarchical Dynamics Synthesis (HDS) mechanism, $\mathcal{HDS}(\cdot;\theta_{\mathcal{HDS}})$, is pivotal to Turb-L1’s strategy for mitigating spectral bias. It operates on the latent representation $\mathbf{z}_{\text{enc}}\in\mathcal{M}_{\omega}$ from the multi-grid encoder $\mathcal{E}$, precisely evolving turbulence dynamics within this manifold. The HDS design leverages insights into the distinct frequency responses of neural operators: Convolutional Neural Networks (CNNs) excel at capturing high-frequency local details, whereas Transformer-like architectures are adept at integrating low-frequency global information [park2022vision, schiaffini2024neural]. Turb-L1 architecturally synergizes these complementary properties for precise, high-fidelity, full-spectrum turbulence evolution.
@@
For synthesizing high-frequency dynamic primitives $\mathbf{h}_{\omega_{H}}$, $\mathcal{T}_{\omega_{H}}$ employs operators highly sensitive to local variations. Convolutional operations, particularly with small receptive field kernels $\kappa_{j}^{(H)}$, inherently function as high-pass filters. Consequently, $\mathcal{T}_{\omega_{H}}$ is instantiated via such convolutional architectures:
$\footnotesize[\mathcal{T}_{\omega_{H}}(\mathbf{z}_{\text{enc}})](\mathbf{x}^{\prime})=\sigma\left(\sum_{j}\int_{\mathcal{N}(\mathbf{x}^{\prime})}\kappa_{j}^{(H)}(\mathbf{x}^{\prime}-\mathbf{y}^{\prime})\mathbf{z}_{\text{enc}}(\mathbf{y}^{\prime})d\mathbf{y}^{\prime}+\mathbf{b}_{j}^{(H)}\right),$ (7)
where $\sigma$ is an activation and $\mathbf{b}_{j}^{(H)}$ a bias. Our implementation, termed HighFrequencyOperator, realizes this with a residual block incorporating a $3\times 3$ depth-wise convolutional positional encoding, a spatial attention mimic composed of a $5\times 5$ depth-wise convolution flanked by $1\times 1$ convolutions, and an MLP block built from $1\times 1$ convolutions.
Conversely, synthesizing low-frequency dynamic primitives $\mathbf{h}_{\omega_{L}}$ is achieved by $\mathcal{T}_{\omega_{L}}$, which targets large-scale structures. The self-attention mechanism (MSA) in Transformers inherently acts as a low-pass filter. $\mathcal{T}_{\omega_{L}}$ is thus instantiable as a Global Adaptive Weighted Aggregation (GAWA) mechanism, rooted in MSA:
$\mathcal{T}_{\omega_{L}}(\mathbf{z}_{\text{enc}})]_{i}=\sum_{k=1}^{N_{h}}\mathbf{W}_{k}^{(O)}\left(\sum_{j=1}^{N_{T}}\alpha_{ij}^{(k)}(\mathbf{W}_{k}^{(V)}\mathbf{z}_{j})\right),$ (8)
where $\alpha_{ij}^{(k)}=\text{softmax}_{j}\left(\frac{(\mathbf{W}_{k}^{(Q)}\mathbf{z}_{i})^{T}(\mathbf{W}_{k}^{(K)}\mathbf{z}_{j})}{\sqrt{d_{k}}}\right)$, $N_{T}$ is the token count, and $\mathbf{W}_{k}^{(\cdot)}$ are learnable head projection matrices. This is implemented as LowFrequencyOperator, a Transformer-style block with standard multi-head self-attention and a linear MLP. Crucially, its residual connections are scaled by learnable parameters initialized to a small value ($10^{-6}$), akin to LayerScale, for stable training.
The concurrently generated dynamic primitives $\{\mathbf{h}_{\omega_{H}},\mathbf{h}_{\omega_{M}},\mathbf{h}_{\omega_{L}}\}$ are subsequently fused by a Multi-scale Dynamic Aggregation (MDA) module, $\mathcal{A}_{\text{MDA}}(\cdot,\cdot,\cdot;\theta_{\mathcal{HDS}}^{(\text{agg})})$. This process yields a unified latent representation $\mathbf{z}_{\text{HDS\_out}}$:
@@
This HDS strategy, with its dedicated processing and intelligent aggregation, forces the model to learn and preserve critical high-frequency details, fundamentally mitigating spectral bias.
@@ 3.5 Mapping to Physical Space
The Multi-Grid Decoder $\mathcal{D}$ receives the latent space features $\mathbf{z}_{\text{HDS\_out}}$. It reconstructs the turbulent vorticity field $\hat{\omega}(\mathbf{x},t_{k+1})$ through a sequence of hierarchical operations. It iteratively applies a Prolongation Operator $\mathcal{P}_{l}$ (upsampling) followed by a learnable Feature Refinement Module $\psi_{l}(\cdot;\theta_{\mathcal{D}}^{(l)})$. In our model, the Prolongation operator is realized using a $3\times 3$ transposed convolution (ConvTranspose2d), again followed by GroupNorm and LeakyReLU. A key feature is a skip-connection that concatenates the upsampled features with the corresponding high-resolution feature map saved from the encoder. This re-introduces fine-scale details lost during downsampling. Finally, an output transformation layer $\mathcal{T}_{\text{out}}$, implemented as a $1\times 1$ convolution, maps the refined features to the final vorticity field. The overall mathematical formulation is:
$\hat{\omega}(\mathbf{x},t_{k+1})=\mathcal{T}_{\text{out}}$ (10)
$\left(\psi_{0}\left(\mathcal{P}_{0}\left(\dots\psi_{L-1}\left(\mathcal{P}_{L-1}\left(\mathbf{z}_{\text{HDS\_out}}\right);\theta_{\mathcal{D}}^{(L-1)}\right)\dots\right);\theta_{\mathcal{D}}^{(0)}\right);\theta_{\mathcal{D}}^{(\text{out})}\right)$
where $\mathbf{z}_{\text{HDS\_out}}$ is the output from the HDS module, $\mathcal{P}_{l}$ denotes the Prolongation operator at stage $l$, and $\psi_{l}$ is the Feature Refinement Module.
3.6 Theoretical Analysis
We theoretically analyze the link between a model’s spectral accuracy and its long-term stability. Consider the true state evolution $\bm{\omega}(t_{k+1})=S(\bm{\omega}(t_{k}))$ and its model approximation $\hat{\bm{\omega}}(t_{k+1})=\mathcal{F}(\hat{\bm{\omega}}(t_{k}))$. The prediction error $\bm{\epsilon}(t_{k})=\hat{\bm{\omega}}(t_{k})-\bm{\omega}^{*}(t_{k})$ evolves approximately as:
@@ 3.6 Theoretical Analysis
Discussion of association with Turb-L1. Theorem 3.1 implies that a smaller $\delta_{\text{HF}}$ is critical for mitigating error growth. The Turb-L1 architecture, with its HDS mechanism, is designed to minimize $\delta_{\text{HF}}$ by explicitly addressing Spectral Bias. This architectural focus on accurately capturing high-frequency dynamics leads to a significant reduction in the model-induced error term in Eq. (13), thereby enhancing long-term stability.
4 Proof of Theorem on High-Frequency Error Control
This section provides a detailed proof for Theorem 3.1.
4.1 Preliminaries: Error Evolution Equation
Let the true state evolution of a dynamical system be governed by $\bm{\omega}(t_{k+1})=S(\bm{\omega}(t_{k}))$. A surrogate model approximates this as $\hat{\bm{\omega}}(t_{k+1})=\mathcal{F}(\hat{\bm{\omega}}(t_{k}))$. The prediction error is $\bm{\epsilon}(t_{k})=\hat{\bm{\omega}}(t_{k})-\bm{\omega}^{*}(t_{k})$. The error at the next step is:
$\bm{\epsilon}(t_{k+1})=\mathcal{F}(\hat{\bm{\omega}}(t_{k}))-S(\bm{\omega}^{*}(t_{k})).$ (14)
A first-order Taylor expansion of $\mathcal{F}(\hat{\bm{\omega}}(t_{k}))$ around $\bm{\omega}^{*}(t_{k})$ gives:
$\mathcal{F}(\hat{\bm{\omega}}(t_{k}))\approx\mathcal{F}(\bm{\omega}^{*}(t_{k}))+\nabla\mathcal{F}|_{\bm{\omega}^{*}(t_{k})}\bm{\epsilon}(t_{k}).$ (15)
Let $\bm{e}_{\text{model}}(t_{k})=\mathcal{F}(\bm{\omega}^{*}(t_{k}))-S(\bm{\omega}^{*}(t_{k}))$ be the single-step model error. Assuming $\nabla\mathcal{F}|_{\bm{\omega}^{*}(t_{k})}\approx\mathbf{A}_{k}=\nabla S|_{\bm{\omega}^{*}(t_{k})}$, the error evolution is:
$\bm{\epsilon}(t_{k+1})\approx\mathbf{A}_{k}\bm{\epsilon}(t_{k})+\bm{e}_{\text{model}}(t_{k}).$ (16)
4.2 Detailed Proof
Proof.
Let $\bm{\epsilon}_{\text{high}}(t_{k})=\mathbf{P}_{\text{high}}\bm{\epsilon}(t_{k})$. Applying $\mathbf{P}_{\text{high}}$ to Eq. (16):
$\bm{\epsilon}_{\text{high}}(t_{k+1})\approx\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}(t_{k})+\mathbf{P}_{\text{high}}\bm{e}_{\text{model}}(t_{k}).$ (17)
Taking the norm and applying the triangle inequality:
$\|\bm{\epsilon}_{\text{high}}(t_{k+1})\|\leq\|\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}(t_{k})\|+\|\mathbf{P}_{\text{high}}\bm{e}_{\text{model}}(t_{k})\|.$ (18)
We assume the amplification of high-frequency error is primarily driven by existing high-frequency components:
$\|\mathbf{P}_{\text{high}}\mathbf{A}_{k}\bm{\epsilon}(t_{k})\|\leq C_{A}\gamma_{\text{high}}\|\bm{\epsilon}_{\text{high}}(t_{k})\|.$ (19)
By the theorem’s assumption, $\|\mathbf{P}_{\text{high}}\bm{e}_{\text{model}}(t_{k})\|\leq\delta_{\text{HF}}$. Substituting these bounds into Eq. (18), we get the recursive inequality:
$\|\bm{\epsilon}_{\text{high}}(t_{k+1})\|\leq C_{A}\gamma_{\text{high}}\|\bm{\epsilon}_{\text{high}}(t_{k})\|+\delta_{\text{HF}}.$ (20)
Let $G=C_{A}\gamma_{\text{high}}$ and $x_{k}=\|\bm{\epsilon}_{\text{high}}(t_{k})\|$. Unrolling the recursion $x_{k+1}\leq Gx_{k}+\delta_{\text{HF}}$ for $M$ steps:
$x_{M}\leq G^{M}x_{0}+\delta_{\text{HF}}\sum_{i=0}^{M-1}G^{i}.$
The geometric series sum is $\frac{G^{M}-1}{G-1}$ for $G\neq 1$. Substituting this back yields the bound:
$\|\bm{\epsilon}_{\text{high}}(t_{M})\|\leq(C_{A}\gamma_{\text{high}})^{M}\|\bm{\epsilon}_{\text{high}}(t_{0})\|+\delta_{\text{HF}}\frac{(C_{A}\gamma_{\text{high}})^{M}-1}{C_{A}\gamma_{\text{high}}-1}.$ (21)
This completes the proof. ∎
4.3 Discussion and Association with Turb-L1
Theorem 3.1 highlights two primary sources of high-frequency error accumulation: amplification of initial error and accumulation of single-step model error. Minimizing the single-step high-frequency error, $\delta_{\text{HF}}$, is critical. The Turb-L1 architecture, with its HDS mechanism, is specifically designed to achieve a smaller $\delta_{\text{HF}}$ by mitigating spectral bias. This explains its superior performance in long-term, high-fidelity turbulence prediction.
@@ 5 Numerical Experiments
TABLE I: Performance comparison on isotropic turbulence datasets measured by $L^{2}$ error at different prediction steps. Lower $L^{2}$ error values indicate better performance.
@@
FNO [li2021fourier] ICLR 2021 0.0267 0.3550 1.7794 2.5634 3.1284 0.0118 0.1384 1.9832
CNO [raonic2023convolutional] NeurIPS 2023 0.0407 1.7774 6.1085 7.7403 11.3015 0.0008 0.1227 1.5676
LSM [wu2023solving] ICML 2023 0.0046 0.3373 3.0698 4.2579 5.1127 0.0017 0.1287 2.0382
NMO [wu2024neural] KDD 2024 0.0018 0.0602 1.7634 1.9832 2.1923 0.0002 0.0043 0.1873
@@
U-Net [ronneberger2015u] MICCAI 2015 0.0182 0.5810 2.8367 3.7935 4.6647 0.0007 0.0296 0.6583
ResNet [he2016deep] CVPR 2016 0.0098 0.4983 1.9823 2.9983 5.8743 0.0025 0.2424 2.3630
ViT [dosovitskiy2021an] ICLR 2021 0.0360 2.2588 5.4557 5.6154 5.6401 0.0074 0.2363 3.8732
DiT [peebles2023scalable] ICCV 2023 0.0038 1.9832 2.7654 5.9862 9.9283 0.0007 0.0283 1.2731
@@
Y ConvLSTM [shi2015convolutional] NIPS 2015 0.0374 0.5643 2.2374 3.9841 4.9381 0.0443 0.0687 1.2384
Y SimVP [tan2022simvp] CVPR 2022 0.0019 0.0621 2.1104 3.8504 5.0405 0.0002 0.0046 0.2231
Y PastNet [wu2024pastnet] MM 2024 0.0128 0.0415 0.5689 1.7574 2.3837 0.0073 0.0348 0.6173
Turb-L1 0.0002 0.0048 0.0807 0.2157 0.6171 0.0001 0.0019 0.1257
@@ 5.1 Experiment Setup
Datasets  We test Turb-L1 on two classic datasets of two-dimensional isotropic turbulence, a canonical problem in fluid dynamics. These datasets are designed to evaluate the model’s capability to capture complex chaotic dynamics under different physical regimes. ✠ Dataset I: Forced Isotropic Turbulence Benchmark. This scenario follows the Navier-Stokes equation benchmark proposed in Li et al. [li2021fourier]. The kinematic viscosity is set to $\nu=10^{-5}$. The initial vorticity field, $\omega_{0}$, is sampled from a zero-mean Gaussian random field with a covariance operator $(-\Delta+\tau^{2}I)^{-\alpha/2}$, where $\Delta$ is the Laplacian operator, $I$ is the identity operator, and $\tau$ and $\alpha$ are parameters controlling the spectral shape. The energy density $E(k)$ exhibits a decay law of $E(k)\sim(k^{2}+\tau^{2})^{-\alpha}$. The flow is driven by a fixed, low-wavenumber external force, and no additional drag effects are considered. ✠ Dataset II: Decaying Isotropic Turbulence (McWilliams, 1984). [mcwilliams1984emergence] This scenario simulates the phenomenon of decaying turbulence. The power spectrum of the initial streamfunction, characterized by $|\hat{\psi}(\mathbf{k})|^{2}\sim k^{-1}(\tau_{0}^{2}+(k/k_{0})^{4})^{-1}$, where $k_{0}$ is the characteristic wavenumber and $\tau_{0}$ is a spectral shape parameter, defines the initial flow field. The initial conditions facilitate a slow energy decay, and the enstrophy density evolves to exhibit features akin to a Kolmogorov energy cascade. The turbulence is unforced and evolves naturally.
Backbones  To comprehensively evaluate the performance of our proposed method, we select three categories of representative deep learning models as baselines for comparison.   Operator Learning Models. These models aim to directly learn mappings between infinite-dimensional function spaces. We choose several mainstream and cutting-edge methods from this field, including: FNO [li2021fourier], CNO [raonic2023convolutional], LSM [wu2023solving], and NMO [wu2024neural].   Classic Computer Vision Backbones. These models demonstrate strong feature extraction capabilities in image processing and computer vision tasks. Specific models include: U-Net [ronneberger2015u], ResNet [he2016deep], ViT [dosovitskiy2021an], and DiT [peebles2023scalable]. Y  Spatiotemporal Prediction Models. These models are specifically designed to process and predict sequential data with temporal and spatial dependencies. Our selected representative models are: ConvLSTM [shi2015convolutional], SimVP [tan2022simvp], and PastNet [wu2024pastnet].
@@
Experiment settings  We train all models in this paper on a server equipped with eight NVIDIA A100 GPUs, each with 40GB of memory. We use the DistributedDataParallel mode from the PyTorch framework for distributed training to accelerate convergence. Model inference is performed on a single NVIDIA A100 GPU. The software environment for our experiments is based on Python 3.8. We use PyTorch version 1.8.1 (CUDA 11.1), TorchVision version 0.9.1. During the training process, we consistently use the following hyperparameter configuration: the batch size is set to 20, the total number of epochs is 500, and the initial learning rate is 0.001. To ensure the reproducibility of our experiments, we fix all random seeds to 42.
Metrics  To comprehensively evaluate our proposed method, we employ five metrics. As follows:
$L^{2}$ Error [heydari2012error]. Measures the $L^{2}$ norm of the error field, which is the difference between the predicted field ($\hat{\omega}(\cdot,t)$) and the ground truth field ($\omega(\cdot,t)$).
$L^{2}\text{ Error}(t)=\|\hat{\omega}(\cdot,t)-\omega(\cdot,t)\|_{L^{2}}$ (22)
$=\sqrt{\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}(\hat{\omega}_{i,j}(t)-\omega_{i,j}(t))^{2}}$
Structural Similarity Index Measure (SSIM) [sampat2009complex]. Assesses the perceptual similarity between two fields (prediction $u$, ground truth $v$) considering local means ($\mu$), variances ($\sigma^{2}$), and covariance ($\sigma_{uv}$).
$\text{SSIM}(u,v)=\frac{(2\mu_{u}\mu_{v}+c_{1})(2\sigma_{uv}+c_{2})}{(\mu_{u}^{2}+\mu_{v}^{2}+c_{1})(\sigma_{u}^{2}+\sigma_{v}^{2}+c_{2})}$ (23)
Relative $L^{2}$ Error [chen2010least]. Normalizes the $L^{2}$ norm of the error field by the $L^{2}$ norm of the ground truth field.
$\text{Relative }L^{2}\text{ Error}(t)=\frac{\|\hat{\omega}(\cdot,t)-\omega(\cdot,t)\|_{L^{2}}}{\|\omega(\cdot,t)\|_{L^{2}}}$ (24)
$=\frac{\sqrt{\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}(\hat{\omega}_{i,j}(t)-\omega_{i,j}(t))^{2}}}{\sqrt{\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}(\omega_{i,j}(t))^{2}}}$
Enstrophy Spectrum (Density), $E_{Z}(k,t)$ [verkley2009energy]. Describes the distribution of enstrophy (mean squared vorticity) across different wavenumbers $k$. Total enstrophy $Z(t)=\int_{0}^{\infty}E_{Z}(k,t)dk$. For 2D fields, it’s computed from the Fourier transform of vorticity $\hat{\omega}(\mathbf{k}^{\prime},t)$ over annular shells in Fourier space, where $A$ is the domain area:
$E_{Z}(k,t)\approx\frac{1}{\Delta k}\sum_{k\leq|\mathbf{k}^{\prime}|<k+\Delta k}\frac{1}{2A}|\hat{\omega}(\mathbf{k}^{\prime},t)|^{2}\cdot(2\pi k^{\prime})$ (25)
Normalized Spectral Error [bazar2016evaluating]. Quantifies the relative difference between the predicted enstrophy spectrum ($E_{Z,\text{pred}}(k,t)$) and the ground truth enstrophy spectrum ($E_{Z,\text{GT}}(k,t)$) at each wavenumber $k$.
$\text{Normalized Spectral Error}(k,t)=\frac{E_{Z,\text{pred}}(k,t)-E_{Z,\text{GT}}(k,t)}{E_{Z,\text{GT}}(k,t)}$ (26)